[Koji, Osamu and Kiwamu]
We found that the ETMX free swinging spectra showed a strange resonant frequencies.
We are going to inspect the suspension today.
In a ideal case the SOS (Small Optic Suspension) is supposed to have the following resonant frequecies.
(Although we didn't carefully identify which corresponds to which)
f_POS ~ 0.98 Hz
f_PITCH ~ 0.66 Hz
f_YAW ~ 0.8 Hz
f_SIDE ~ 0.99 Hz
However ETMX showed the following resonant frequencies.
f_POS ~ 0.91 Hz
f_PITCH ~ 0.7 Hz
f_YAW ~ 0.93 Hz
f_SIDE ~ 1.0 Hz
Especially f_YAW looks pretty high. Also the others are not at the right frequencies.
So we are suspicious that something wrong is happening on the ETMX suspension.